### v1

Abstract
Time-dependent partial differential equations are ubiquitous in physics-based modeling, but they remain computationally intensive in many-query scenarios, such as real-time forecasting, optimal control, and uncertainty quantification. Reduced-order modeling (ROM) addresses these challenges by constructing a low-dimensional surrogate model but relies on a fixed discretization, which limits flexibility across varying meshes during evaluation. Operator learning approaches, such as neural operators, offer an alternative by parameterizing mappings between infinite-dimensional function spaces, enabling adaptation to data across different resolutions. Whereas ROM provides rigorous numerical error estimates, neural operator learning largely focuses on discretization convergence and invariance without quantifying the error between the infinite-dimensional and the discretized operators. This work introduces the reduced-order neural operator modeling (RONOM) framework, which bridges concepts from ROM and operator learning. We establish a discretization error bound analogous to those in ROM, and get insights into RONOM’s discretization convergence and discretization robustness. Moreover, two numerical examples are presented that compare RONOM to existing neural operators for solving partial differential equations. The results demonstrate that RONOM using standard vector-to-vector neural networks achieves comparable performance in input generalization and superior performance in both spatial super-resolution and discretization robustness, while also offering novel insights into temporal super-resolution scenarios.
keywords:
Reduced-order modeling, operator learning, error estimates, machine learning, partial differential equations
{MSCcodes}
65D15, 65D40, 68W25, 65M99, 68T20, 68T07
1 Introduction
Time-dependent partial differential equations (PDEs) are fundamental for physics-based modeling for a wide range of real-world systems [38]. However, the complexity and scale of these systems often make model evaluations computationally expensive or even intractable. This is especially problematic in many-query scenarios, such as real-time forecasting [29, 5, 31], optimal design and control [5, 36], and uncertainty quantification [45, 48], which typically require multiple evaluations of the parametric model with varying physical parameters, initial conditions, or boundary conditions. One of the state-of-the-art methods to mitigate this problem is reduced-order modeling (ROM) [37, 33]. ROM leverages the coherent structure over space, time, or even the parameter space to construct a low-dimensional manifold. The low-dimensional representation of the system resolves the efficiency issue of model evaluation without significantly compromising the accuracy of the model [37]. One common way to construct the manifold is proper orthogonal decomposition (POD) [40]. However, the subspace spanned by the linear bases of POD may suffer from slowly decaying Kolmogorov n-width for transport problems, advection-dominated PDEs, or non-affine parametrizations of domain geometries and physical variables [37]. A variety of nonlinear dimensionality reduction methods have been proposed to address this issue. In particular, deep learning approaches based on autoencoder neural networks have gained considerable attention for their flexibility and strong performance [22, 16, 15].
One common limitation of those linear and nonlinear dimensionality reduction methods is that they inherit the spatial discretization from solution snapshots generated by traditional numerical methods, e.g., finite difference and finite element methods (FEM). For instance, the low-dimensional POD representation is formulated in a finite-dimensional subspace spanned by those snapshots, and the bases constructed via low-rank approximation are dependent on the specific discretization of the snapshots [37]. Similarly, the nonlinear manifold in nonlinear model reduction also relies on solutions at a pre-defined discretization. Consequently, these models struggle to generalize across varying discretizations, which is often required in many-query scenarios, e.g., when locally refining grids to resolve fine-scale features.
One way to mitigate the impact of discretization, but not in the context of ROM, is operator learning [6]. Operator learning approximates a mapping between functions. It can be applied to approximate the mapping from an initial condition or source term to the solution of the considered PDE. Such an approximation at the continuous level avoids the explicit dependency on discretization and allows zero-shot super-resolution. Methods such as the random feature model [28], DeepONet [26], Fourier neural operator (FNO) [23], and graph neural operator (GNO) [24] have demonstrated their applicability to different PDE problems. Within the broad class of neural operator (NO) methods, several incorporate latent structures. Some use a linear combination of basis functions to map to functions [4, 17, 19]. Others leverage implicit neural representations for more flexible and powerful nonlinear maps [42, 49, 50, 12].
Whereas neural operators are commonly presented as function-to-function maps, in practice, they work on discretized input functions. The learned operators are usually only discretization-convergent, meaning that the discretized operators converge to a true infinite-dimensional operator as the input resolution is refined. However, as it is only a convergence property, it does not guarantee robustness to discretization changes. The recent ReNO framework [2] addresses this limitation by designing operators with a form of equivalence between the input function and its discrete samples. In particular, the CNO [39] assumes bandlimited functions and leverages the fact that bandlimited functions satisfy the Nyquist–Shannon theorem, which ensures equivalence between functions and their point samples. Consequently, any input over the Nyquist rate can be projected back to the same rate of training before being passed to the operator, yielding equal performance when tested on higher-resolution scenarios. But on the other hand, such a setting also limits its applicability.
While the properties of neural operators, such as discretization convergence and discretization invariance, are widely explored in existing literature, it often lacks a thorough analysis of how well the discretized operators approximate their underlying infinite-dimensional counterparts. To the best of our knowledge, only Lanthaler et al. [21] investigate this error by analyzing the discretization errors made by FNOs. In contrast, several ROM methods are equipped with rigorous error analysis, guaranteeing the accuracy of the reduced system’s numerical construction and approximation [37]. This is closely related to discretization convergence and discretization invariance. If the error estimates guarantee that refining the numerical solution brings it closer to the true solution, then the numerical solver is discretization convergent.
This work proposes the reduced-order neural operator modeling (RONOM) framework that connects operator learning and ROM. The general structure of RONOM is illustrated in Figure 1. An architecture based on standard vector-to-vector neural networks is adopted, inspired by the kernel method of Batlle et al. [3]. We demonstrate that appropriate lifting and sampling turn these standard neural network structures, which also appear in ROM, into a neural operator with competitive performance. We prove a discretization error bound for RONOM and evaluate its performance in terms of the three desired properties illustrated in Figure 2, namely generalization to unseen inputs, super-resolution, and discretization robustness. Moreover, it is important to note that the latent code in this work is not necessarily a reduced representation in the traditional ROM sense. Since functions are considered and their discrete resolution may vary, it is more accurate to view the latent code as a finite-dimensional representation (projection) of the function at arbitrary resolution. The main contributions of this work are summarized as follows:
Figure 1: RONOM is a modeling framework that combines the strengths of neural operators and reduced order modeling. It begins by mapping the initial condition into a latent space. From this initial latent representation, a latent trajectory is obtained over time. The full trajectory of functions is recovered by decoding the latent codes at each time instance.
Figure 2: Robustness and superresolution through RONOM. Traditional finite-dimensional ROM methods for PDE surrogate modeling can generalize to new inputs, provided these input functions use the same fixed discretization as in training (black arrow). In contrast, RONOM can also enhance the resolution beyond the training resolution (blue arrow), and it remains applicable and robust when inputs are given at different discretizations than those used in training (green arrows).
  • Modeling: a discretization-convergent neural operator is proposed that connects traditional ROM to recent developments in neural operator learning.
  • Analysis: discretization error bounds are established for the neural operator, unifying numerical error analysis in ROMs with the discretization convergence concept of neural operators.
  • Numerics: numerical experiments provide insight into the temporal super-resolution of operator learning and demonstrate that, i) comparable input generalization performance is achieved compared to the FNO, CNO, and DeepONet, and ii) spatial super-resolution and discretization robustness are achieved in cases where FNOs and CNOs fail.
2 The reduced-order neural operator model
This section introduces the details of RONOM. The specific time-dependent PDEs that RONOM is designed to solve are first described. The model architecture, including the encoder, decoder, and neural ordinary differential equation (ODE) components, is subsequently detailed. Discretization error estimates and Lipschitz continuity results are also presented. These results are crucial for providing RONOM’s error estimates in Section 3.2.
2.1 Problem statement
Consider a time-dependent PDE on a bounded domain $\Omega\subset\mathbb{R}^{d}$,
(1) $\frac{\partial}{\partial t}u(\bm{x},t)+\mathcal{N}(u,\bm{x})=0,$ $\hskip 99.58464pt(\bm{x},t)\in\Omega\times[0,T]$
$\mathcal{B}u(\bm{x},t)=g(\bm{x}),$ $\hskip 99.58464pt(\bm{x},t)\in\partial\Omega\times[0,T]$
$u(\bm{x},0)=u_{0}(\bm{x}),$ $\hskip 99.58464pt\bm{x}\in\Omega.$
where $u$ denotes the solution of the PDE; $\mathcal{N}$ is a general differential operator and $\mathcal{B}$ is the boundary condition operator defined on the domain boundary $\partial\Omega$. In this work, we aim to learn the nonlinear solution operator $\mathcal{K}\colon(u_{0},t)\mapsto u(\cdot,t)$ based on supervised training data.
2.2 Overview of RONOM
RONOM aims to approximate the solution operator $\mathcal{K}$ by an operator $\mathcal{K}^{\dagger}$ consisting of three maps,
$\mathcal{K}(u_{0},t)\approx\mathcal{K}^{\dagger}(u_{0},t)\coloneqq(D\circ\mathcal{F}_{t}\circ E)(u_{0}).$
The map $E\colon u_{0}\mapsto\mathbf{z}_{0}\in\mathbb{R}^{d_{z}}$ is the encoder that maps the initial condition to a finite-dimensional latent code of dimension $d_{z}$. Such a latent space allows us to explore and march over time, which is analogous to the reduced system of the reduced basis method [37, 29]. Our model realizes the time marching in latent space through the flow operator $\mathcal{F}_{t}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{z}}$. More specifically, $\mathcal{F}_{t}(\mathbf{z}_{0})=\mathbf{z}(t)$ with $\mathbf{z}(t)$ the solution of a first-order ODE system,
(2) $\frac{\mathrm{d}\mathbf{z}(t)}{\mathrm{d}t}=\mathbf{v}(\mathbf{z}(t),t),\quad\mathbf{z}(0)=\mathbf{z}_{0}=E(u_{0}),$
where $\mathbf{v}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{z}}$ determines the latent dynamics. Finally, the decoder $D:\mathbb{R}^{d_{z}}\to L_{2}(\Omega)$ recovers the approximated solution from the latent code $\mathbf{z}(t)$.
Each component of $\mathcal{K}^{\dagger}$ is characterized by neural networks. Given the available training data $\{u_{i0},\{u_{ij}\coloneq u_{i}(t_{j})\}_{j=0}^{N_{t}-1}\}_{i=1}^{N_{s}}$, the learning of the solution operator $\mathcal{K}^{\dagger}$ relies on minimizing the objective function,
$\min_{\bm{\theta}}$ $\left(\frac{1}{N_{s}N_{t}}\sum_{i=1}^{N_{s}}\left(\sum_{j=0}^{N_{t}-1}\norm{u_{ij}-\widehat{u}_{i}(t_{j})}_{L_{2}(\Omega)}^{2}+\mu_{1}\norm{u_{ij}-\widehat{u}_{ij}}_{L_{2}(\Omega)}^{2}\right)\right.$
$\left.\qquad+\mu_{2}\norm{\bm{\mathrm{z}}_{i0}}_{2}^{2}+\mu_{3}R_{1}\left(\{\bm{\mathrm{z}}_{i}(t_{j})\}_{j=0}^{N_{t}-1}\right)+\mu_{4}R_{2}\left(\{\bm{\mathrm{z}}_{i}(t_{j})\}_{j=0}^{N_{t}-1}\right)\right)+\mu_{5}\norm{\bm{\theta}}_{2}^{2}$
s.t. $\mathbf{z}_{ij}=E\left(u_{ij}\right),\quad\widehat{u}_{ij}=D(\bm{\mathrm{z}}_{ij}),$
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{\mathrm{z}}_{i}(t)=\mathbf{v}(\bm{\mathrm{z}}_{i}(t),t),\quad\bm{\mathrm{z}}_{i}(0)=\bm{\mathrm{z}}_{i0},\quad\widehat{u}_{i}(t)=D(\bm{\mathrm{z}}_{i}(t)),$
where $\bm{\theta}$ represents the set of all trainable parameters; $\mu_{i}\in\mathbb{R}$; $R_{1}$ is a velocity field magnitude regularizer and $R_{2}$ is a smoothness regularizer,
$R_{1}\left(\{\bm{\mathrm{z}}_{i}(t_{j})\}_{j=0}^{N_{t}-1}\right)=$ $\sum_{j=0}^{N_{t}-2}\norm{\frac{\bm{\mathrm{z}}_{i}(t_{j+1})-\bm{\mathrm{z}}_{i}(t_{j})}{\delta_{t}}}_{2}^{2},$
$R_{2}\left(\{\bm{\mathrm{z}}_{i}(t_{j})\}_{j=0}^{N_{t}-1}\right)=$ $\sum_{j=1}^{N_{t}-2}\norm{\frac{\bm{\mathrm{z}}_{i}(t_{j+1})-2\bm{\mathrm{z}}_{i}(t_{j})+\bm{\mathrm{z}}_{i}(t_{j-1})}{\delta_{t}^{2}}}_{2}^{2},$
where $t_{i+1}-t_{i}=\delta_{t}$, $\forall i\in\{0,\ldots,N_{t}-2\}$.
2.3 Encoder
To ensure the discretization convergence and discretization robustness of the encoder, inspiration is drawn from the optimal recovery approach presented in Batlle et al. [3] and the frame sequence approach employed in the CNO [39]. Both approaches map the input samples to a full function, followed by the application of a predefined linear operator to obtain a finite-dimensional vector representation.
Our encoder first uses a regularized $L_{2}$ projection $\mathcal{P}_{\mathcal{V}}^{\lambda}$ to project the input function to a finite-dimensional subspace $\mathcal{V}:=\operatorname{span}\{\phi_{1},\phi_{2},\cdots,\phi_{N_{b}}\}\subset L_{2}(\Omega)$, where $\{\phi_{i}\}_{i=1}^{N_{b}}$ are linearly independent. In particular, the projection $\mathcal{P}_{\mathcal{V}}^{\lambda}$ is written as,
(3) $\mathcal{P}_{\mathcal{V}}^{\lambda}f:=\operatorname*{argmin}_{u\in\mathcal{V}}\int_{\Omega}|u(\bm{x})-f(\bm{x})|^{2}\mathrm{d}\bm{x}+\lambda\norm{\mathcal{L}u}_{R}^{2},$
where $\mathcal{L}$ is some (invertible) linear operator on $\mathcal{V}$ and $\norm{\cdot}_{R}$ a norm coming from an inner product |$\innerproduct{\cdot}{\cdot}_{R}$. To compute this projection for discretized input functions in the form of grids, meshes, or point clouds, a discretize-then-optimize approach is applied, which is discussed in Section 2.3.1. After the projection, we apply a measurement operator $M\colon\mathcal{V}\to\mathbb{R}^{d_{m}}$, for instance by sampling the projection at specific points or by mapping it to the basis coefficients [4, 17, 1]. Subsequently, the resulting vector is processed as in a DeepONet. In particular, with a neural network $\mathcal{E}_{\varphi}\colon\mathbb{R}^{d_{m}}\to\mathbb{R}^{d_{z}}$, the full (infinite-dimensional) encoder is given by:
(4) $E(u_{0}):=(\mathcal{E}_{\varphi}\circ M\circ\mathcal{P}_{\mathcal{V}}^{\lambda})(u_{0}).$
For a discussion on how this encoder connects to the frame-sequence approach used by the CNO, see LABEL:app:connection_encoder_and_CNO. Note that it is important to ensure that the measurement operator $M$ is injective when restricted to the subspace $\mathcal{V}$. This injectivity guarantees the equivalence between the measurement values and the (projected) functions in $\mathcal{V}$. To ensure injectivity, the matrix $\mathrm{M}_{\phi}\in\mathbb{R}^{d_{m}\times N_{b}}$ with entries $(\mathrm{M}_{\phi})_{\cdot i}=M\phi_{i}$, must have full rank $N_{b}$, which implies the condition $d_{m}\geq N_{b}$. Therefore, the basis functions spanning $\mathcal{V}$ should be chosen such that $N_{b}\leq d_{m}$, and the measurement operator $M$ must be designed such that the matrix $\mathrm{M}_{\phi}$ has rank $N_{b}$.
2.3.1 Discretized projection
To apply the encoder in Equation (4) for data in the form of a mesh or point cloud, the $\mathcal{P}_{\mathcal{V}}^{\lambda}$ mapping is discretized using a discretize-then-optimize approach. The discretized projection solves a regularized least-squares problem that maps the input $\{\bm{x}_{i},f(\bm{x}_{i})\}_{i=1}^{n}=\{\bm{x}_{i},f_{i}\}_{i=1}^{n}$ to a function
(5) $\widehat{\mathcal{P}}_{\mathcal{V}}^{\lambda}(\{\bm{x}_{i},f_{i}\}_{i=1}^{n})\coloneqq\operatorname*{argmin}_{u\in\mathcal{V}}\sum_{i=1}^{n}w_{i}|u(\bm{x}_{i})-f_{i}|^{2}+\lambda\norm{\mathcal{L}u}_{R}^{2},$
where $w_{i}$ are weighting constants used to approximate the data fidelity integral in Equation (3). The convergence of the solution of Equation (5) to the solution of Equation (3) will be discussed in the next section. The following theorem shows that the regularized projection can be interpreted as a specific regularized instance of the kernel approach in Batlle et al. [3].
Theorem 2.1.
Assuming $\mathcal{L}$ is invertible on $\mathcal{V}=\operatorname{span}\{\phi_{1},\phi_{2},\cdots,\phi_{N_{b}}\}$ and define the $\mathcal{H}_{K}$ inner product as $\innerproduct{f}{g}_{\mathcal{H}_{K}}=\innerproduct{\mathcal{L}f}{\mathcal{L}g}_{R}$. Moreover, let $\tilde{\phi}_{i}$ be an orthonormal basis of the Hilbert Space $(\mathcal{V},\innerproduct{\cdot}{\cdot}_{\mathcal{H}_{K}})$. Define the kernel $K\colon\Omega\times\Omega\to\mathbb{R}$ as $K(\bm{x},\bm{y})=\sum_{i=1}^{N_{b}}\tilde{\phi}_{i}(\bm{x})\tilde{\phi_{i}}(\bm{y})$. If $w_{i}=1$, optimization problem (5) is equivalent to the soft-constraint version of the optimal recovery problem:
(6) $\min_{u\in\mathcal{H}_{K}}\sum_{i=1}^{n}|u(\bm{x}_{i})-f_{i}|^{2}+\lambda\norm{u}_{H_{K}}^{2},$
with $\mathcal{H}_{K}=\mathcal{V}$ the reproducing kernel Hilbert space corresponding to $K$.
Proof 2.2.
Given the definition of $(\mathcal{H}_{K},\innerproduct{\cdot}{\cdot}_{\mathcal{H}_{K}})$ and $w_{i}=1$, we get the equivalence of Equations (5) and (6). To show $K$ is the reproducing kernel of $\mathcal{H}_{K}$, note that
$\innerproduct{K(\cdot,\bm{x})}{\tilde{\phi}_{j}}_{\mathcal{H}_{K}}=\sum_{i=1}^{N_{b}}\tilde{\phi}_{i}(\bm{x})\innerproduct{\tilde{\phi_{i}}}{\tilde{\phi}_{j}}_{\mathcal{H}_{K}}=\sum_{i=1}^{N_{b}}\tilde{\phi}_{i}(\bm{x})\delta_{ij}=\tilde{\phi}_{j}(\bm{x}).$
Therefore for any $f=\sum_{i=1}^{N_{b}}a_{i}\tilde{\phi}_{i}\in\mathcal{H}_{K}$, $a_{i}\in\mathbb{R}$, the following holds,
$\innerproduct{K(\cdot,\bm{x})}{f}_{\mathcal{H}_{K}}=\sum_{i=1}^{N_{b}}a_{i}\innerproduct{K(\cdot,\bm{x})}{\tilde{\phi}_{i}}_{\mathcal{H}_{K}}=\sum_{i=1}^{N_{b}}a_{i}\tilde{\phi}_{i}(\bm{x})=f(\bm{x}),$
which confirms that $K$ serves as the reproducing kernel of $\mathcal{H}_{K}=\mathcal{V}$ with respect to the inner product |$\innerproduct{\cdot}{\cdot}_{\mathcal{H}_{K}}$.
2.3.2 Error estimate projection’s basis coefficients
To get an encoder error estimate in Section 3.2, it is necessary to bound the difference between the solutions to (3) and (5). By expressing $u(\bm{x})=\sum^{N_{b}}_{i=1}\alpha_{i}\phi_{i}(\bm{x})=\bm{\phi}(\bm{x})^{\top}\bm{\alpha}$, the problems (3) and (5) can be reformulated as follows,
(7) $\min_{\bm{\alpha}\in\mathbb{R}^{N_{b}}}\bm{\alpha}^{\top}\Phi\bm{\alpha}-2\bm{\alpha}^{\top}\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}+\norm{f}_{L_{2}(\Omega)}^{2}+\lambda\bm{\alpha}^{\top}\mathrm{L}\bm{\alpha}.$
and
(8) $\min_{\bm{\alpha}\in\mathbb{R}^{N_{b}}}\bm{\alpha}^{\top}\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}\bm{\alpha}-2\bm{\alpha}^{\top}\tilde{\Phi}\mathrm{D}_{w}\bm{f}+\bm{f}^{\top}\mathrm{D}_{w}\bm{f}+\lambda\bm{\alpha}^{\top}\mathrm{L}\bm{\alpha}$
respectively, where $(\mathrm{L})_{ij}=\innerproduct{\mathcal{L}\phi_{i}}{\mathcal{L}\phi_{j}}_{R}$, $\tilde{\Phi}_{ij}=\phi_{i}(\bm{x}_{j})$, $(\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)})_{i}=\innerproduct{f}{\phi_{i}}_{L_{2}(\Omega)}$, $(\mathrm{D}_{w})_{ij}=w_{i}\delta_{ij}$, and $\bm{f}=[f_{1},\cdots,f_{n}]^{\top}$. Hence, the difference between the solutions is determined by the difference between the basis coefficients. As shown in the next theorem, this is mainly controlled by how well $\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}$ approximates $\Phi$ and how well $\tilde{\Phi}\mathrm{D}_{w}\bm{f}$ approximates $\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}$,
Theorem 2.3.
Assume $\bm{\alpha}^{\dagger}$ solves Equation (7) and $\bm{\alpha}$ Equation (8). Assuming $\mathrm{L}$ is invertible, the difference between them can be bounded by:
$\norm{\bm{\alpha}-\bm{\alpha}^{\dagger}}_{2}$ $\leq\frac{\norm{\mathrm{L}^{-1}}_{2}}{\lambda}\Big{(}\norm{\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}-\tilde{\Phi}\mathrm{D}_{w}\bm{f}}_{2}$
$+\norm{(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}-\Phi)}_{2}\norm{(\Phi+\lambda\mathrm{L})^{-1}}_{2}\norm{\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}}_{2}\Big{)}.$
Proof 2.4.
Putting the gradient with respect to $\bm{\alpha}$ to zero in problems (7) and (8) gives us $(\Phi+\lambda\mathrm{L})\bm{\alpha}^{\dagger}=\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}$ and $(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})\bm{\alpha}=\tilde{\Phi}\mathrm{D}_{w}\bm{f}$. Moreover,
$(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})(\bm{\alpha}^{\dagger}-\bm{\alpha})=(\Phi+\lambda\mathrm{L})(\bm{\alpha}^{\dagger})-(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})\bm{\alpha}+(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}-\Phi)\bm{\alpha}^{\dagger}$
By the optimality conditions above, this becomes
(9) $\left(\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}-\tilde{\Phi}\mathrm{D}_{w}\bm{f}\right)+(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}-\Phi)(\Phi+\lambda\mathrm{L})^{-1}\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}.$
To obtain a bound on the difference $\bm{\alpha}-\bm{\alpha}^{\dagger}$, the triangle inequality is applied as a first step,
$\norm{\bm{\alpha}-\bm{\alpha}^{\dagger}}_{2}$ $=\norm{(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})^{-1}(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})(\bm{\alpha}-\bm{\alpha}^{\dagger})}_{2}$
$\leq\norm{(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})^{-1}}_{2}\norm{(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})(\bm{\alpha}-\bm{\alpha}^{\dagger})}_{2}$
Finally, using $\norm{(\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{\top}+\lambda\mathrm{L})^{-1}}_{2}\leq\frac{\norm{\mathrm{L}^{-1}}_{2}}{\lambda}$ and applying the triangle inequality on (9) completes the proof.
As stated in the theorem, it suffices to quantify how accurately $\tilde{\Phi}\mathrm{D}_{w}\tilde{\Phi}^{T}$ approximates $\Phi$, as well as how well $\tilde{\Phi}\mathrm{D}_{w}\bm{f}$ approximates $\innerproduct{f}{\bm{\phi}}_{L_{2}(\Omega)}$. To this end, note that
$\left(\tilde{\Phi}D_{w}\tilde{\Phi}^{\top}\right)_{ij}=\sum_{k=1}^{N_{b}}w_{k}\phi_{i}(\bm{x}_{k})\phi_{j}(\bm{x}_{k}),\quad\left(\tilde{\Phi}\mathrm{D}_{w}\bm{f}\right)_{i}=\sum_{k=1}^{N_{b}}w_{k}\phi_{i}(\bm{x}_{k})f(\bm{x}_{k}).$
Therefore, it suffices to establish error estimates for $L_{2}$ inner products.
2.3.3 Integral discretization error estimates
Given a mesh, the standard approach to approximate the (inner product) integrals is numerical integration, also known as numerical quadrature. To understand and assess the impact of quadrature errors on the approximation of the $L_{2}$ inner products, an error bound is presented below. The detailed proof is provided in Appendix A.
Theorem 2.5.
Assume that the compact domain $\Omega$ is approximated by a mesh $\Omega^{N}\subseteq\Omega$ consisting of $N$ mesh elements $\Omega_{k}$ with nonempty interior, i.e., $\operatorname{int}(\Omega_{k})$: $\Omega^{N}=\bigcup_{k=1}^{N}\Omega_{k}$ and $\operatorname{int}(\Omega_{k})\cap\operatorname{int}(\Omega_{\tilde{k}})=\emptyset$ for $k\neq\tilde{k}$. Assume either:
  • $f\in C^{p+1}(\Omega)$ and $\{w_{i,k},\bm{x}_{i,k}\}_{i=1}^{m_{k}}$ is a quadrature rule of order $p$, meaning $\int_{\Omega_{k}}q(\bm{x})\mathrm{d}\bm{x}=\sum_{i=1}^{m}w_{i,k}q(\bm{x}_{i,k})$ for all $p$-th order polynomials $q$ on $\Omega_{k}$.
  • $f\in\operatorname{TV}(\Omega)$, $w_{i,k}=|\Omega_{k}|\widetilde{w}_{i,k}$ with $\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}=1$, and $\bm{x}_{i,k}\in\Omega_{k}$.
Then $\left|\int_{\Omega}f(\bm{x})\mathrm{d}\bm{x}-\sum_{k=1}^{N}\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right|$ is bounded by
$|\Omega\Delta\Omega^{N}|\norm{f}_{L_{\infty}(\Omega)}+\begin{cases}ch^{p+1}|\Omega^{N}|\sup_{\begin{subarray}{c}\bm{x}\in\Omega\\ \sum_{i=1}^{d}\gamma_{i}=p+1\end{subarray}}\left|\partial^{\gamma}f(\bm{x})\right|,&\quad f\in C^{p+1}(\Omega)\\ h\operatorname{TV}(f),&\quad f\in\operatorname{TV}(\Omega),\end{cases}$
where $c\in\mathbb{R}$ only depends on $p$, $\partial^{\gamma}f(\bm{x}):=\frac{\partial^{|\gamma|}}{\partial x_{1}^{\gamma_{1}}\cdots\partial x_{d}^{\gamma_{d}}}f(\bm{x})$ with a multi-index notation $\gamma=(\gamma_{1},\gamma_{2},\cdots,\gamma_{d})$, $h_{\Omega_{k}}\coloneqq\sup_{\bm{x}_{1},\bm{x}_{2}\in\Omega_{k}}\norm{\bm{x}_{1}-\bm{x}_{2}}_{2}\leq h$, $\Omega\Delta\Omega^{N}:=(\Omega\setminus\Omega^{N})\cup(\Omega^{N}\setminus\Omega)=\Omega\ \setminus\Omega^{N}$ since $\Omega^{N}\subseteq\Omega$, and $|\Omega\Delta\Omega^{N}|$ is its area.
Remark 2.6.
The estimate above highlights several important aspects. First, in regions where the function is sufficiently smooth, rapid convergence is observed. In contrast, regions with discontinuities exhibit significantly slower convergence due to the inherent difficulty of approximating abrupt changes.
This issue could arise, for example, when approximating a function defined on a domain $\Omega$ using basis functions supported on a larger domain $\widetilde{\Omega}$, where $\Omega\subset\widetilde{\Omega}$. In particular, Theorem 2.3 shows that part of the projection error stems from approximating the inner products $\innerproduct{f}{\phi_{i}}_{L_{2}(\widetilde{\Omega})}$. If $f$ is only defined on $\Omega$, this inner product is typically computed by extending $f$ to zero outside $\Omega$, which might introduce a discontinuity at the boundary of $\Omega$. This discontinuity makes accurate integration more difficult near the boundary of $\Omega$.
Finally, approximating the domain itself introduces an additional source of error.
The previous error estimate requires a mesh. To deal with point clouds, one can consider Monte Carlo sampling to approximate the integral, assume the samples are sampled according to some specific distribution, and obtain a probabilistic bound with an error of $\mathcal{O}(1/\sqrt{N})$ [41]. To remove the randomness assumption and get a non-probabilistic bound, one can employ quasi-Monte Carlo schemes. Instead of randomly sampling and approximating the integral in that way, one uses a sum $\frac{1}{N}\sum_{i=1}^{N}f(\bm{x}_{i})$ with deterministic $\bm{x}_{i}\in\Omega$ to approximate $\int_{\Omega}f(\bm{x})\mathrm{d}\bm{x}$. In this case, one can employ Koksma-Hlawka type inequalities to bound the approximation error [30, 20, 32]. The general form of such inequalities bounds the error as a product of two terms:
  • a discrepancy term that measures how well the sampled points are approximating the given distribution or measure, and
  • a measure of variation of the integrand.
In particular, Theorem 4.3 of [32] proves a Koksma-Hlawka inequality when integrating over general compact domains. For completeness, the complete statement of the theorem is provided in LABEL:SM:koksma-hlawka_ineq.
2.4 Neural ODE flow operator
After encoding the input into an initial latent code $\mathbf{z}(0)=\mathbf{z}_{0}=E(u_{0})$, the neural ODE in Equation (2) returns a latent trajectory $\mathbf{z}(t)$. Any numerical solver can be used; however, it only yields values at the points $\{t_{i}\}_{i=0}^{N_{\delta_{t}}-1}$ that are used to solve the differential equation. It is common to employ an interpolation method to obtain latent vectors at intermediate time points, which is necessary for evaluating the neural operator at arbitrary values of $t$.
Given a time discretization $\bm{t}:=\{t_{i}\}_{i=0}^{N_{\delta_{t}}-1}\subset[0,T]$ with $N_{\delta_{t}}$ time instances, a numerical solver $\Psi(\mathbf{z};\bm{t})$ can be used to approximate the discrete solution at times $t_{i}$,
$\Psi(\mathbf{z};\bm{t})=\{\widehat{\mathbf{z}}(t_{i})\}_{i=0}^{N_{\delta_{t}}-1},\quad\text{with }\widehat{\mathbf{z}}(t_{0})=\mathbf{z},$
where $\widehat{\mathbf{z}}(t_{i})$ denotes the approximated discrete solution of the solver at time instance $t_{i}$ and $\mathbf{z}\in\mathbb{R}^{d_{z}}$ denotes the initial condition. To recover a continuous representation over time from the discrete solutions, Hermite spline interpolations can be applied. In the scalar case, Hermite interpolation constructs a polynomial $s(t)$ of degree $12p-$ 1 that approximates a general smooth function $f:[0,h]\to\mathbb{R}$ over a small interval, with the following conditions on the polynomial $s$,
$\frac{\mathrm{d}^{i}}{\mathrm{d}x^{i}}s(0)=\frac{\mathrm{d}^{i}}{\mathrm{d}x^{i}}f(0),\quad\frac{\mathrm{d}^{i}}{\mathrm{d}x^{i}}s(h)=\frac{\mathrm{d}^{i}}{\mathrm{d}x^{i}}f(h),\quad\quad\text{for }i=0,\dots,p-1,$
Since $\widehat{\mathbf{z}}(t_{i})\in\mathbb{R}^{d_{z}}$, Hermite interpolation can be applied component-wise on each interval $[t_{i},t_{i+1}]$. For the $i$-th interval $[t_{i},t_{i+1}]$ and $j$th coordinate $j=1,\ldots,d_{z}$, a univariate Hermite interpolant $s_{ij}(t)$ can be constructed. The full interpolated value for $t\in[t_{i},t_{i+1}]$ ($i\in\{0,\dots,N_{\delta_{t}}-2\}$) can be written as $\bm{s}(t)\coloneqq\bm{s}_{i}(t)$ with $\bm{s}_{i}(t)=[s_{i1}(t),s_{i2}(t),\cdots,s_{id_{z}}(t)]^{\top}$.
For demonstration, cubic interpolation is assumed here, namely $p=2$. The cubic interpolation uses $\{\widehat{\mathbf{z}}(t_{i})\}_{i=0}^{N_{\delta_{t}}-1}$ and the exact time derivatives $\mathbf{v}(\widehat{\mathbf{z}}(t_{i}),t)$. Combining the numerical solver with the interpolation scheme yields a time-continuous approximation ^$\widehat{\mathcal{F}}_{t}(\mathbf{z};\bm{t})$ of the exact flow map $\mathcal{F}_{t}(\mathbf{z})$. In particular, denoting the interpolation as $\bm{s}(t;\Psi(\mathbf{z};\bm{t}),\bm{t})$ to highlight its dependency on the time-discretization and the numerical solver, the time-continuous approximation is
(10) ^$\widehat{\mathcal{F}}_{t}(\mathbf{z};\bm{t})\coloneqq\bm{s}(t;\Psi(\mathbf{z};\bm{t}),\bm{t}).$
The choice of numerical solver and the order of the spline interpolation method directly determine the error between ^$\widehat{\mathcal{F}}_{t}$ and $\mathcal{F}_{t}$. While this holds for any $p\geq 1$ (as shown in LABEL:SM:p_unequal_to_2_case_error_estimate_NODE), we demonstrate the case when $p=2$.
Theorem 2.7.
Assume we have a numerical ODE solver $\Psi$ with global error order $\mathcal{O}(\delta_{t}^{q})$ and a time discretization $\bm{t}:=\{t_{i}\}_{i=0}^{N_{\delta_{t}}-1}$ satisfying $\sup_{i\in\{1,\ldots,N_{\delta_{t}}-1\}}|t_{i}-t_{i-1}|\leq\delta_{t}$. Furthermore, assume $\mathbf{v}$ is $L_{v}$-Lipschitz in $\mathbf{z}$ and that for $j=1,\ldots,d_{z}$ the functions $(\mathbf{z},t)\mapsto(R_{3}(\mathbf{z},t))_{j}$ are in $L^{\infty}(Z\times\mathbb{R})$ for $R_{0}(\mathbf{z},t)\coloneqq\mathbf{v}(\mathbf{z},t)$ and $R_{k}$ defined as:
$R_{k}(\mathbf{z},t)\coloneqq\left(\sum_{j=0}^{d_{z}}\sum_{l=0}^{k-1}{k-1\choose l}\left(\frac{\partial}{\partial\mathrm{z}_{j}}R_{k-1-l}(\mathbf{z},t)\right)(R_{l}(\mathbf{z},t))_{j}\right)+\frac{\partial}{\partial t}R_{k-1}(\mathbf{z},t).$
Then, when using cubic Hermite interpolation in Equation (10), the following error estimate applies,
$\norm{\mathcal{F}_{t}(\mathbf{z})-\widehat{\mathcal{F}}_{t}(\mathbf{z};\bm{t})}_{2}=\mathcal{O}\left(\delta_{t}^{\min(4,q)}\right).$
Proof 2.8.
Take a time $t\in[t_{i},t_{i+1}]$. Then $\mathcal{F}_{t}(\mathbf{z}):=\mathbf{z}(t)$ and ^$\widehat{\mathcal{F}}_{t}(\mathbf{z};\bm{t})=\bm{s}_{i}(t)$. Define the numerical approximation $\widehat{\mathbf{z}}(\bm{t}):=\{\widehat{\mathbf{z}}(t_{i})\}_{i=0}^{N_{\delta_{t}}-1}$ to Equation (2). Moreover, let $\tilde{\bm{s}}_{i}$ be defined analogously to $\bm{s}_{i}$, but constructed by interpolating the ground truth values $\mathbf{z}(t_{i})$ and $\mathbf{z}(t_{i+1})$. Then for $t\in[t_{i},t_{i+1}]$ it follows that,
(11) $\begin{split}\norm{\mathbf{z}(t)-\bm{s}_{i}(t)}_{2}&\leq\norm{\mathbf{z}(t)-\tilde{\bm{s}}_{i}(t)}_{2}+\norm{\tilde{\bm{s}}_{i}(t)-\bm{s}_{i}(t)}_{2}\\ &\leq\frac{\delta_{t}^{4}}{2^{4}(4)!}\left({\sum_{j=1}^{d_{z}}\norm{\frac{\mathrm{d}^{4}}{\mathrm{d}t^{4}}\mathrm{z}_{j}}_{L^{\infty}(t_{i},t_{i+1})}^{2}}\right)^{\frac{1}{2}}+\norm{\tilde{\bm{s}}_{i}(t)-\bm{s}_{i}(t)}_{2},\end{split}$
where the final inequality follows from standard Hermite interpolation bounds [43, 10]. To show the existence and boundedness of the fourth-order derivative, first observe that,
$\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\mathrm{z}_{i}=\frac{\mathrm{d}^{k-1}}{\mathrm{d}\mathrm{t}^{k-1}}v_{i}(\mathbf{z}(t),t)=\frac{\mathrm{d}^{k-2}}{\mathrm{d}t^{k-2}}\left(\sum_{j=0}^{d_{z}}\frac{\partial v_{i}(\mathbf{z}(t),t)}{\partial\mathrm{z}_{j}}\frac{\mathrm{d}\mathrm{z}_{j}}{\mathrm{d}t}+\frac{\partial v_{i}(\mathbf{z}(t),t)}{\partial t}\right),$
which can be rewritten using the general Leibniz rule to,
$\left(\sum_{j=0}^{d_{z}}\sum_{l=0}^{k-2}{k-2\choose l}\left(\frac{\partial}{\partial\mathrm{z}_{j}}\frac{\mathrm{d}^{k-2-l}}{\mathrm{d}t^{k-2-l}}(v_{i}(\mathbf{z}(t),t))\right)\frac{\mathrm{d}^{l}}{\mathrm{d}t^{l}}v_{j}(\mathbf{z}(t),t)\right)+\frac{\partial}{\partial t}\frac{\mathrm{d}^{k-2}}{\mathrm{d}t^{k-2}}(v_{i}(\mathbf{z}(t),t).$
Hence, $\frac{\mathrm{d}^{k}}{\mathrm{d}t^{k}}\mathrm{z}_{j}=(R_{k-1}(\mathbf{z}(t),t))_{j}$, which implies $\norm{\frac{\mathrm{d}^{4}}{\mathrm{d}t^{4}}\mathrm{z}_{j}}_{L^{\infty}(t_{i},t_{i+1})}^{2}<\infty$ as $(\mathbf{z},t)\mapsto(R_{3}(\mathbf{z},t))_{j}$ is a function in $L^{\infty}(Z\times\mathbb{R})$.
By defining $\delta_{i}:=(t_{i+1}-t_{i})$, $h_{00}(x)\coloneqq(2x^{3}-3x^{2}+1)$, $h_{10}(x)\coloneqq(x^{3}-2x^{2}+x)$, $h_{01}(x)\coloneqq(-2x^{3}+3x^{2})$, $h_{11}(x)\coloneqq(x^{3}-x^{2})$, and $\tilde{t}\coloneqq\frac{t-t_{i}}{\delta_{i}}$, the cubic spline interpolation is given by $\tilde{\bm{s}}_{i}(t)=h_{00}(\tilde{t})\mathbf{z}(t_{i})+h_{01}(\tilde{t})\mathbf{z}(t_{i+1})+h_{10}(\tilde{t})\delta_{i}\frac{\mathrm{d}\mathbf{z}(t_{i})}{\mathrm{d}t}+h_{11}(\tilde{t})\delta_{i}\frac{\mathrm{d}\mathbf{z}(t_{i+1})}{\mathrm{d}t}$. With this formulation for the spline interpolation and defining $\mathbf{e}(t)\coloneqq\mathbf{z}(t)-\widehat{\mathbf{z}}(t)$, the interpolation error on each interval can be written as,
$\begin{split}\norm{\tilde{\bm{s}}_{i}(t)-\bm{s}_{i}(t)}_{2}&=\norm{\sum_{k=0}^{1}h_{0k}(\tilde{t})\mathbf{e}(t_{i+k})+h_{1k}(\tilde{t})\delta_{i}\left(\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{e}(t_{i+k})\right)}_{2}\\ &\leq\sum_{k=0}^{1}|h_{0k}(\tilde{t})|\norm{\mathbf{e}(t_{i+k})}_{2}+|h_{1k}(\tilde{t})|\delta_{i}\norm{\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{e}(t_{i+k})}_{2}.\end{split}$
For $j=i,i+1$, $\norm{\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{e}(t_{j})}_{2}=\norm{\frac{\mathrm{d}\mathbf{z}(t_{j})}{\mathrm{d}t}-\frac{\mathrm{d}\widehat{\mathbf{z}}(t_{j})}{\mathrm{d}t}}_{2}=\norm{\mathbf{v}(\mathbf{z}(t_{j}),t_{j})-\mathbf{v}(\widehat{\mathbf{z}}(t_{j}),t_{j})}_{2}$, and hence $\norm{\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{e}(t_{j})}_{2}\leq L_{v}\norm{\mathbf{z}(t_{j})-\widehat{\mathbf{z}}(t_{j})}=L_{v}\norm{\mathbf{e}(t_{j})}$ by Lipschitz-continuity of $\mathbf{v}$. Combining this result with the preceding inequality yields,
$\begin{split}\norm{\tilde{\bm{s}}_{i}(t)-\bm{s}_{i}(t)}_{2}&\leq(|h_{00}(\tilde{t})|+L_{v}|h_{10}(\tilde{t})|\delta_{i})\norm{\mathbf{z}(t_{i})-\widehat{\mathbf{z}}(t_{i})}_{2}\\ &+(|h_{01}(\tilde{t})|+L_{v}|h_{11}(\tilde{t})|\delta_{i})\norm{\mathbf{z}(t_{i+1})-\widehat{\mathbf{z}}(t_{i+1})}_{2}\\ &\leq\left(|h_{00}(\tilde{t})|+|h_{01}(\tilde{t})|+L_{v}(|h_{10}(\tilde{t})|+|h_{11}(\tilde{t})|)\delta_{t}\right)C\delta_{t}^{q},\end{split}$
where the last inequality holds as the numerical solver is of the order $q$ and $\delta_{i}\leq\delta_{t}$. By integrating this result with the inequality shown in (11) and the boundedness of $h_{kl}$ on $[0,1]$, we get an error of order $\delta_{t}^{\min(4,q)}$, confirming the claim.
Remark 2.9.
The $\mathbf{v}$ in Equation (2) is parameterized by a neural network, which is trained jointly with the encoder and decoder. In the reduced basis method, the encoder and decoder are constructed before solving the reduced system via an optimization problem. A similar optimization approach is used in the Neural Galerkin method [8] and in evolutionary deep neural networks [18]. These approaches first fit a neural network to the initial condition. Its parameters are subsequently progressively updated along directions that minimize the PDE residual over time. Similarly, given a current estimate $\mathbf{z}(t)$ and noting that $\frac{\partial}{\partial t}D(\mathbf{z}(t))(\bm{x})=\innerproduct{\nabla_{z}D(\mathbf{z})(\bm{x})}{\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{z}(t)}$, we can also choose the velocity vector field that corresponds to the best projection of $\mathcal{N}(D(\mathbf{z}(t)),\cdot)$ onto the tangent space of our latent manifold,
$\min_{\dot{\mathbf{z}}:=\frac{\mathrm{d}}{\mathrm{d}t}\bm{\mathbf{z}}(t)}\int_{\Omega}\norm{\innerproduct{\nabla_{z}D(\mathbf{z})(\bm{x})}{\dot{\mathbf{z}}}-\mathcal{N}(D(\mathbf{z}(t)),\bm{x})}^{2}\mathrm{d}\bm{x}.$
For a given $\bm{x}$, let $J_{D}(\mathbf{z},\bm{x})$ denote the Jacobian of $D(\mathbf{z})(\bm{x})$ with respect to $\mathbf{z}$. The following matrix equation gives the solution to this problem,
$\left(\int_{\Omega}J_{D}(\mathbf{z}(t),\bm{x})^{T}J_{D}(\mathbf{z}(t),\bm{x})\mathrm{d}\bm{x}\right)\dot{\mathbf{z}}=\left(\int_{\Omega}J_{D}(\mathbf{z},\bm{x})\mathcal{N}(D(\mathbf{z}(t)),\bm{x})\mathrm{d}\bm{x}\right).$
To apply Theorem 2.7, it is necessary to show that $(\mathbf{z},t)\mapsto(R_{3}(\mathbf{z},t))_{j}$ is in $L^{\infty}(Z\times\mathbb{R})$ for $j=1,\ldots,d_{z}$ and that the solutions to the above system of equations are Lipschitz in $\mathbf{z}(t)$. The latter property parallels the result established in Theorem 2.3. Overall, under appropriate assumptions on the decoder, Theorem 2.7 remains valid.
2.5 Decoder
The decoder maps latent vectors $\mathbf{z}(t)$ to functions. A typical example is the DeepONet, where $\mathbf{z}(t)$ feeds into the branch net and spatial coordinates into the basis functions. To align with standard ROM methods, a DeepONet-style decoder is employed, utilizing fixed, non-learnable basis functions. This structure closely relates to the POD-DeepONet [27], which chooses the basis a-priori via POD.
In particular, our decoder is based on the optimal recovery method from Batlle et al. [3]. Let $\{\bm{x}_{i}\}_{i=1}^{N_{c}}\subset\Omega$ be a set of distinct spatial locations, and let $K\colon\Omega\times\Omega\to\mathbb{R}$ be a positive definite kernel associated to a reproducing kernel Hilbert space $\mathcal{H}_{K}$ with norm $\|\cdot\|_{K}$. The regularized version of the optimal recovery problem seeks the function $u$ that solves:
$\min_{u\in\mathcal{H}_{K}}$ $\sum_{i=1}^{N_{c}}|u(\bm{x}_{i})-y_{i}|^{2}+\eta\|u\|_{K}^{2},$
where $\eta\rightarrow\infty$ corresponds to the optimal recovery problem, which finds $u$ of minimal RKHS norm that interpolates the given data. Let the coefficients $\{a_{i}\}_{i=1}^{N_{c}}$ originate from the linear system $K_{\eta}\bm{a}=\mathbf{y}$, where $K_{\eta}\coloneqq K(\mathbf{X},\mathbf{X})+\eta I$, $\bm{a}=[a_{1},\dots,a_{N_{c}}]^{\top}$ and $(\mathbf{X},\mathbf{y})=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{c}}$. The solution to the regularized problem can be written as,
$u\left(\bm{x}\,\big{|}\,\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{c}}\right)=\sum_{i=1}^{N_{c}}a_{i}K(\bm{x},\bm{x}_{i})=K(\bm{x},\mathbf{X})K_{\eta}^{-1}\mathbf{y}.$
When the solution operator to the PDE is trained, the values $y_{i}$ are unknown. Consequently, a map $D_{d}\colon Z\to\mathbb{R}^{N_{c}}$ is defined from the latent space to the predicted values $y_{i}$. The complete form of the decoder is then expressed as,
(12) $D(\mathbf{z})(\bm{x})=\sum_{i=1}^{N_{b}}\left(K_{\eta}^{-1}D_{d}(\mathbf{z})\right)_{i}K(\bm{x},\bm{x}_{i})=K(\bm{x},\mathbf{X})K_{\eta}^{-1}D_{d}(\mathbf{z}).$
An important feature of this construction is that $D_{d}$ can be any latent-to-vector map used in the ROM literature. The decoder outputs a function via upsampling with a kernel, and its Lipschitz continuity can be analyzed straightforwardly for any norm on functions.
Theorem 2.10.
Let $\norm{\cdot}_{\mathcal{U}}$ be a norm on functions and $D_{d}$ be $L_{D_{d}}$-Lipschitz. Then $D$ in Equation (12) is $L_{D}$-Lipschitz in $\norm{\cdot}_{\mathcal{U}}$ where $L_{D}:=L_{D_{d}}C_{F}\sqrt{\norm{K_{\eta}^{-T}AK_{\eta}^{-1}}_{2}}$ and $A$ is defined by:
$A_{ij}=\begin{cases}\innerproduct{K(\cdot,\bm{x}_{i})}{K(\cdot,\bm{x}_{j})}_{\mathcal{U}},\quad\norm{\cdot}_{\mathcal{U}}\text{ comes from an inner product}\\ K(\bm{x}_{i},\bm{x}_{j}),\quad\,\,\qquad\qquad\text{otherwise}\end{cases}$
and $C_{F}=1$ when $\norm{\cdot}_{\mathcal{U}}$ comes from an inner product and depends on $\norm{\cdot}_{\mathcal{U}}$ and $K$ otherwise.
Proof 2.11.
For the case that an inner product does not induce the norm, we can define $\innerproduct{\sum_{i=1}^{N_{c}}a_{i}K(\cdot,\bm{x}_{i})}{\sum_{i=1}^{N_{c}}b_{j}K(\cdot,\bm{x}_{j})}_{\mathcal{H}_{K}}:=\sum_{i,j=1}^{N_{c}}a_{i}K(\bm{x}_{i},\bm{x}_{j})b_{j}$ as the inner product on our space $\operatorname{span}(K(\cdot,\bm{x}_{i})\mid i=1,\ldots,N_{c})$. By equivalence of norms on finite-dimensional spaces, $\norm{D(\mathbf{z}_{1})-D(\mathbf{z}_{2})}_{\mathcal{U}}\leq\widetilde{C}_{\mathcal{U}}\norm{D(\mathbf{z}_{1})-D(\mathbf{z}_{2})}_{\mathcal{H}_{K}}$ for some constant $\widetilde{C}_{\mathcal{U}}$. When considering the $\mathcal{H}_{K}$ inner product or when considering a norm $\norm{\cdot}_{\mathcal{U}}$ that comes from an inner product, Equation (12) gives us that $\norm{D(\mathbf{z}_{1})-D(\mathbf{z}_{2})}^{2}$ can be rewritten to,
$\innerproduct{\sum_{i=1}^{N_{c}}(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))_{i}K(\cdot,\bm{x}_{i})}{\sum_{j=1}^{N_{c}}(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))_{j}K(\cdot,\bm{x}_{j})}$
$=\sum_{i,j=1}^{N_{c}}(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))_{i}\innerproduct{K(\cdot,\bm{x}_{i})}{K(\cdot,\bm{x}_{j})}(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))_{j},$
where the inner product is |$\innerproduct{\cdot}{\cdot}_{\mathcal{U}}$ when $\norm{\cdot}_{\mathcal{U}}$ comes from an inner product or |$\innerproduct{\cdot}{\cdot}_{\mathcal{H}_{K}}$ otherwise. Let $C_{F}=1$ when $\norm{\cdot}_{\mathcal{U}}$ comes from an inner product and let $C_{F}=\widetilde{C}_{\mathcal{U}}$ otherwise. Combining the previous observations with the definition of the matrix $A$, it leads to
$\norm{D(\mathbf{z}_{1})-D(\mathbf{z}_{2})}_{\mathcal{U}}^{2}$ $\leq C_{F}^{2}(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))^{\top}A(K_{\eta}^{-1}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2})))$
$=C_{F}^{2}(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2}))^{\top}\left(K_{\eta}^{-\top}AK_{\eta}^{-1}\right)(D_{d}(\mathbf{z}_{1})-D_{d}(\mathbf{z}_{2}))$
$\leq C_{F}^{2}\norm{K_{\eta}^{-\top}AK_{\eta}^{-1}}_{2}\norm{D(\mathbf{z}_{1})-D(\mathbf{z}_{2})}_{2}^{2}$
$\leq L_{D_{d}}^{2}C_{F}^{2}\norm{K_{\eta}^{-\top}AK_{\eta}^{-1}}_{2}\norm{\mathbf{z}_{1}-\mathbf{z}_{2}}_{2}^{2}=L_{D}^{2}\norm{\mathbf{z}_{1}-\mathbf{z}_{2}}_{2}^{2}.$
3 RONOM error estimates
Discretization plays an essential role in both operator learning and ROM. In the neural operator literature, discretization convergence and discretization invariance characterize whether discretizations of a learned infinite-dimensional operator converge to this infinite-dimensional operator. In the ROM literature, similar questions arise when analyzing the discretization error of the reduced system. We present a theorem that, for the first time, bridges these questions in neural operator learning and ROM. Before stating this result, we provide a general a-posteriori error estimate for the discrepancy between our infinite-dimensional reconstruction and the ground truth solution of the PDE.
3.1 An a-posteriori error estimate
The following theorem is closely related to Proposition 1 from Farenga et al. [15]. Whereas their result is formulated for ODEs, our theorem is established in the context of PDEs. In addition, our error estimate is expressed solely in terms of quantities that can be computed from the model, in contrast to theirs, which also depends on the unknown ground truth solution.
Theorem 3.1.
Assume the strong solution $u(t)$ to the PDE in Equation (1) exists and let $\hat{u}(t):=D(\mathbf{z}(t))$ be our approximation with $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{z}(t)=\mathbf{v}(\mathbf{z}(t),t)$ with $\mathbf{z}(0)=\mathbf{z}_{0}$. Furthermore, suppose a Hilbert space $\mathcal{U}$ is given such that $u(t),\hat{u}(t)\in\mathcal{U}$ and $\mathcal{N}(u,\cdot)\in\mathcal{U}\,\,\forall u\in\mathcal{U}$. Assuming the nonlinear operator $\mathcal{N}$ is $L_{\mathcal{N}}$-Lipschitz with respect to $\norm{\cdot}_{\mathcal{U}}$, the error $e(t):=\norm{u(t)-\hat{u}(t)}_{\mathcal{U}}$ satisfies,
$e(t)\leq\mathrm{e}^{L_{\mathcal{N}}t}\left(e(0)+\int_{0}^{t}\left(\norm{\mathcal{N}(\hat{u}(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}\right)\mathrm{d}t\right).$
Proof 3.2.
Note that $\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}e(t)^{2}\right)=\innerproduct{u(t)-\hat{u}(t)}{\frac{\partial}{\partial t}(u(t)-\hat{u}(t))}_{\mathcal{U}}$ and:
$\innerproduct{u(t)-\hat{u}(t)}{\frac{\partial}{\partial t}(u(t)-\hat{u}(t))}_{\mathcal{U}}=\innerproduct{u(t)-\hat{u}(t)}{\mathcal{N}(u(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}$
$=\innerproduct{u(t)-\hat{u}(t)}{\mathcal{N}(u(t),\cdot)-\mathcal{N}(\hat{u}(t),\cdot)}_{\mathcal{U}}+\innerproduct{u(t)-\hat{u}(t)}{\mathcal{N}(\hat{u}(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}$
$\leq L_{\mathcal{N}}\norm{u(t)-\hat{u}(t)}_{\mathcal{U}}^{2}+\norm{\mathcal{N}(\hat{u}(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}\norm{u(t)-\hat{u}(t)}_{\mathcal{U}}$
$=L_{\mathcal{N}}e(t)^{2}+\norm{\mathcal{N}(\hat{u}(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}e(t).$
The inequality for the first term arises from the fact that the differential operator $\mathcal{N}$ is $L_{\mathcal{N}}$-Lipschitz, and the second term follows from the Cauchy–Schwarz inequality. Using $e(t)\frac{\mathrm{d}}{\mathrm{d}t}e(t)=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}e(t)^{2}\right)$ and dividing both sides by $e(t)$ then yields,
$\frac{\mathrm{d}}{\mathrm{d}t}e(t)\leq L_{\mathcal{N}}e(t)+\norm{\mathcal{N}(\hat{u}(t),\cdot)-\frac{\partial}{\partial t}\hat{u}(t)}_{\mathcal{U}}.$
Using Gronwall’s inequality yields the desired inequality.
The inequality highlights two key aspects. It shows that errors in encoding the initial condition propagate through $e(0)$. If the initial condition is reconstructed perfectly, the remaining error stems from $\frac{\partial}{\partial t}\hat{u}(t)$ not precisely following the evolution dictated by the PDE.
3.2 Discretization error estimate
As mentioned in the introduction, neural operator frameworks are frequently developed without a thorough analysis of how accurately their discretized implementations approximate the underlying infinite-dimensional operators. In contrast, ROM techniques, especially reduced basis methods, are supported by rigorous error analyses, which provide theoretical guidance on the accuracy of the numerical construction and approximation of the reduced system for time-dependent problems.
In the theorem below, we unify these perspectives by providing an error estimate for the full RONOM pipeline, which also establishes discretization convergence.
Theorem 3.3 (Discretization error estimate RONOM).
Assume that $\mathcal{E}_{\varphi}$ in the encoder is $L_{\mathcal{E}_{\varphi}}$-Lipschitz, the decoder $D\colon Z\to\mathcal{U}\subset L_{2}(\Omega)$ is $L_{D}$-Lipschitz for some Banach Space $\mathcal{U}$ (e.g., see Theorem 2.10), and the assumptions of Theorem 2.7 are satisfied. Let $\widehat{\mathcal{K}}^{\dagger}(U_{0},t)\coloneqq(D\circ\widehat{\mathcal{F}}_{t}\circ\widehat{E})(U_{0})$ approximate $\mathcal{K}^{\dagger}(u_{0},t)(\bm{x}):=(D\circ\mathcal{F}_{t}\circ E)(u_{0})(\bm{x})$, where $U_{0}\coloneqq\{\bm{x}_{i},u_{0i}\}_{i=1}^{n}$, ^$\widehat{\mathcal{F}}_{t}(\mathbf{z};\bm{t})$ is given in Equation (10) and the discrete encoder $\widehat{E}$ results from approximating $\mathcal{P}_{\mathcal{V}}^{\lambda}$ using $\widehat{\mathcal{P}}_{\mathcal{V}}^{\lambda}$, i.e.,
$\widehat{E}(U_{0})\coloneqq(\mathcal{E}_{\varphi}\circ M\circ\widehat{\mathcal{P}}_{\mathcal{V}}^{\lambda})(U_{0}).$
Then with $q$ the order of the numerical ODE solver, $L_{v}$ the Lipschitz constant of $\mathbf{v}$, time discretization $\bm{t}:=\{t_{i}\}_{i=0}^{N_{\delta_{t}}-1}$ satisfying $\sup_{i=1,\ldots,N_{\delta_{t}}-1}|t_{i}-t_{i-1}|\leq\delta_{t}$, $\mathrm{M}_{\phi}$ the matrix with columns $M\phi_{i}$, and $L(t)\coloneqq L_{D}e^{L_{v}t}\norm{\mathrm{M}_{\phi}}_{2}$, the following holds,
$\norm{\mathcal{K}^{\dagger}(u_{0},t)-\widehat{\mathcal{K}}^{\dagger}(U_{0},t)}_{\mathcal{U}}\leq L(t)e_{a}(u_{0},U_{0})+L_{D}\mathcal{O}\left(\delta_{t}^{\min(4,q)}\right),$
where $e_{a}(u_{0},U_{0})$ denotes the bound from Theorem 2.3,
$e_{a}(u_{0},U_{0})$ $\coloneqq\frac{\norm{\mathrm{L}^{-1}}_{2}}{\lambda}\Big{(}\norm{\innerproduct{u_{0}}{\bm{\phi}}_{L_{2}(\Omega)}-\tilde{\Phi}D_{w}\bm{u_{0}}}_{2}$
$+\norm{(\tilde{\Phi}D_{w}\tilde{\Phi}^{T}-\Phi)}_{2}\norm{(\Phi+\lambda\mathrm{L})^{-1}}_{2}\norm{\innerproduct{u_{0}}{\bm{\phi}}_{L_{2}(\Omega)}}_{2}\Big{)}.$
Proof 3.4.
An error estimate for the discretized encoder is first established. This result is then combined with previously derived estimates to obtain a comprehensive error bound for the overall framework.
Error estimate between $E$ and $\widehat{E}$
Let $a^{\dagger}(u_{0})$ and $a(U_{0})$ be the coefficients such that $\mathcal{P}_{\mathcal{V}}^{\lambda}(u_{0})=\sum_{i=1}^{N_{b}}a^{\dagger}_{i}(u_{0})\phi_{i}$ and $\widehat{\mathcal{P}}_{\mathcal{V}}^{\lambda}(U_{0})=\sum_{i=1}^{N_{b}}a_{i}(U_{0})\phi_{i}$, respectively. Owing to linearity of $M$ and $L_{\mathcal{E}_{\varphi}}$ Lipschitzness of $\mathcal{E}_{\varphi}$, we obtain the following estimate:
$\norm{E(u_{0})-\widehat{E}(U_{0})}_{2}$ $=\norm{\mathcal{E}_{\varphi}\circ M\circ\mathcal{P}_{\mathcal{V}}^{\lambda}(u_{0})-\mathcal{E}_{\varphi}\circ M\circ\widehat{\mathcal{P}}_{\mathcal{V}}^{\lambda}(U_{0})}_{2}$
$\leq L_{\mathcal{E}_{\varphi}}\norm{\left(\sum_{i=1}^{N_{b}}(a_{i}^{\dagger}(u_{0})-a_{i}(U_{0}))M\phi_{i}\right)}_{2}$
$=L_{\mathcal{E}_{\varphi}}\norm{\mathrm{M}_{\phi}(a^{\dagger}(u_{0})-a(U_{0}))}_{2}$
$\leq L_{\mathcal{E}_{\varphi}}\norm{\mathrm{M}_{\phi}}_{2}\norm{a^{\dagger}(u_{0})-a(U_{0})}_{2}.$
Combining all error estimates
Let $\mathbf{z}(t)$ and $\tilde{\mathbf{z}}(t)$ denote the solution of the neural ODE given initial conditions $\mathbf{z}$ and $\mathbf{z}+\delta_{z}$, respectively. Then [46, Theorem 2.8] yields $\norm{\mathcal{F}_{t}(\mathbf{z})-\mathcal{F}_{t}(\mathbf{z}+\delta_{z})}_{2}\leq\norm{\delta_{z}}_{2}e^{L_{v}t}$. Utilizing this stability estimate, we obtain,
$\norm{\mathcal{K}^{\dagger}(u_{0},t)-\widehat{\mathcal{K}}^{\dagger}(U_{0},t)}_{\mathcal{U}}$ $=\norm{D(\mathcal{F}_{t}(E(u_{0})))-D(\widehat{\mathcal{F}}_{t}(\widehat{E}(U_{0});\bm{t})}_{\mathcal{U}}$
$\leq L_{D}\norm{\mathcal{F}_{t}(E(u_{0}))-\widehat{\mathcal{F}}_{t}(\widehat{E}(U_{0});\bm{t})}_{2}$
$\leq L_{D}\left(\norm{\mathcal{F}_{t}(E(u_{0}))-\mathcal{F}_{t}(\widehat{E}(U_{0}))}_{2}\right.$
$\left.+\norm{\mathcal{F}_{t}(\widehat{E}(U_{0}))-\widehat{\mathcal{F}}_{t}(\widehat{E}(U_{0});\bm{t})}_{2}\right)$
$\leq L_{D}\left(e^{L_{v}t}\norm{E(u_{0})-\widehat{E}(U_{0})}_{2}+\mathcal{O}\left(\delta_{t}^{\min(4,q)}\right)\right)$
$\leq L_{D}\left(e^{L_{v}t}L_{\mathcal{E}_{\varphi}}\norm{\mathrm{M}_{\phi}}_{2}\norm{a^{\dagger}(u_{0})-a(U_{0})}_{2}+\mathcal{O}\left(\delta_{t}^{\min(4,q)}\right)\right),$
where the second-to-last inequality follows from the stability estimate and Theorem 2.7, and the last inequality follows from the error estimate between $E$ and $\widehat{E}$. The full inequality now follows from Theorem 2.3.
Remark 3.5.
The error estimate for the encoder depends on three main factors: the conditioning of the matrix $\Phi+\lambda\mathrm{L}$, and the approximation of the inner products in $\innerproduct{u_{0}}{\bm{\phi}}_{L_{2}(\Omega)}$ and in $\Phi$. As shown in Section 2.3.3, these inner product approximation errors are influenced by domain approximation and numerical quadrature. Error estimates similar to those of our encoder also arise in FEM, where variational crimes account for errors due to, for instance, domain approximation, inexact matching of the boundary conditions, and numerical quadrature [44, 7].
Beyond the effects of domain approximation and numerical quadrature, the conditioning of $\Phi+\lambda\mathrm{L}$ and the FEM assembly matrix plays a critical role. In FEM, refining the mesh typically deteriorates the conditioning of the assembly matrix, which in turn can amplify quadrature errors when computing the right-hand side of the system. While FEM analysis focuses on how the condition number evolves with increasing mesh resolution, our approach employs a fixed global basis, rendering $\Phi$ independent of the mesh. Such construction requires only a single condition number analysis, irrespective of the discretization of the input function.
4 Numerical experiments
This section compares RONOM, FNO, CNO, and DeepONet based on the three properties in Figure 2, namely input generalization, super-resolution, and discretization robustness. The comparisons are demonstrated through two time-dependent PDE examples, the one-dimensional Burgers’ equation and the two-dimensional wave equation.
To address time-dependency with the FNO and CNO, the time variable is added as a constant input feature at each spatial grid point. The trunk network takes both space and time as input for DeepONet. Since the data introduced in the next section is given on grids, RONOM’s encoder combines convolutional neural networks (CNNs), a grid-based sampling, and regularized projections. These projections use Gaussian basis functions centered at grid points and incorporate Sobolev norms, orthonormality penalties, and boundary constraints. The latent representation matches the spatial resolution, augmented with four channels; for example, it is $33\times 4$ for Burgers’ equation and $33\times 33\times 4$ for the wave equation. The decoder maps latent codes back to the grid using a CNN and applies a Gaussian kernel for optimal recovery. More details on RONOM’s design choices can be found in LABEL:SM:ronom_arch_details.
4.1 Data generation
4.1.1 Burgers’ equation
Consider a one-dimensional viscous Burgers’ equation with periodic boundary conditions,
$\begin{cases}\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},&(x,t)\in(0,1)\times(0,1]\\ u(0,t)=u(1,t),&t\in(0,1]\\ \frac{\partial u}{\partial t}(0,t)=\frac{\partial u}{\partial t}(1,t),&t\in(0,1]\\ u(x,0)=u_{0}(x),&x\in(0,1)\end{cases}$
where $\nu=0.01$ is the viscosity coefficient. The initial condition $u_{0}$ is generated from a Gaussian random field with mean zero, and as the covariance kernel, the operator $\sqrt{2}(250^{2})(-\Delta+25^{2}I)^{-2}$ is considered with periodic boundary conditions. A total of 1000 solution trajectories are generated using Chebfun [11] on a spatial grid of size $1025$, with $101$ time points uniformly spaced between 00 and $1$. To prepare the training data, the spatial and temporal grids are subsampled by factors of $32$ and $10$, respectively. This yields a spatial resolution of $33$, with time instances $[0.0,0.1,\ldots,1.0]$ in training. Among $1000$ solutions, $800$ of them are used for training and validation, while the remaining $200$ are for testing.
4.1.2 Wave equation
Consider a two-dimensional wave equation with wave propagation speed $c=0.3$, Dirichlet boundary conditions, initial condition $f$, and zero initial velocity:
$\begin{dcases}\frac{\partial^{2}}{\partial t^{2}}u=c^{2}\Delta u,&(\bm{x},t)\in(0,1)^{2}\times(0,1]\\ u(\bm{x},t)=0,&(\bm{x},t)\in\partial[0,1]^{2}\times[0,1]\\ u(\bm{x},0)=f(\bm{x}),&\bm{x}\in[0,1]^{2}\\ u_{t}(\bm{x},0)=0,&\bm{x}\in[0,1]^{2},\end{dcases}$
where $f(\bm{x})=\sum_{i,j=1}^{K}a_{ij}(\pi^{2}(i^{2}+j^{2}))^{-r/2}\sin(\pi ix_{1})\sin(\pi jx_{2})$ denotes initial condition and the analytical solution of the PDE can be written as $u(\bm{x},t)=\sum_{i,j=1}^{K}a_{ij}(\pi^{2}(i^{2}+j^{2}))^{-r/2}\sin(\pi ix_{1})\sin(\pi jx_{2})\mathrm{cos}(c\pi t\sqrt{i^{2}+j^{2}}),$ where $K=24$, $r=3$, $a_{11}=0$, and $a_{ij}\sim\mathcal{N}(0,125^{2})$. In this example, $1200$ solutions are generated based on a spatial grid of size $129\times 129$, with $101$ time points uniformly spaced between 00 and $1$. For training, the spatial grid is subsampled by a factor of 4 in each dimension, and the temporal grid is subsampled by a factor of 10. This yields a training resolution of $33\times 3$ 33, with time instances $[0.0,0.1,\ldots,1.0]$. Among the entire dataset, $1000$ of them are used for training and validation, while the remaining $200$ is used for testing.
4.2 Results on input generalization and super-resolution
In this section, the results concerning input generalization and super-resolution are presented. To demonstrate those two properties, the input discretization used during evaluation is kept consistent with the training resolution.
For input generalization, models are evaluated by mapping inputs and querying outputs at the same resolution as used during training. In the super-resolution setting, the input remains at the training resolution, while the output is generated at a higher spatial resolution. Specifically, this means generating solutions on the full grid with $1025$ points for the Burgers’ equation and $129\times 129$ grid points for the wave equation. To assess the super-resolution performance regarding time, outputs are evaluated at both the training and full time resolutions. Table 1 summarizes the quantitative results based on the root mean square error (RMSE).
Table 1: Evaluating input generalization and super-resolution. The average RMSE of FNO, CNO, DeepONet, and RONOM predictions on the test dataset. The values are of the order $10^{-2}$, and the best values are highlighted in bold. As FNO requires the input and output spatial grids to match, it can not be used for the spatial super-resolution experiment.
Models Input generalization Super-resolution
Time Space + Time
Burgers Wave Burgers Wave Burgers Wave
FNO 0.075 0.083 0.245 0.086 $\times$ $\times$
CNO 0.31 0.098 0.482 0.301 2.459 1.750
DeepONet 1.24 1.385 1.459 1.344 1.353 1.380
RONOM 0.36 0.093 0.503 0.645 0.508 0.663
4.2.1 Input generalization
Table 1 shows that the FNO method outperforms both CNO and RONOM on the Burgers’ equation, while CNO and RONOM achieve similar RMSE around 0.3. All three methods perform comparably on the wave equation case. In contrast, DeepONet performs significantly worse in both scenarios.
To further illustrate the differences among the methods, visualizations are shown in Figure 3 for the Burgers’ equation and in Figure 4 for the wave equation. Figure 3 shows that the FNO demonstrates much less reconstruction error, and it outperforms CNO, RONOM, and DeepONet on the Burgers’ equation. Moreover, both CNO and RONOM appear to struggle with accurately reconstructing the initial condition. This is also observed in the wave equation predictions presented in Figure 4. The discrepancy in the initial condition reconstructions likely stems from the architectural differences since CNO and RONOM both rely on local convolutional layers, while the FNO allows copying the full initial condition via global Fourier layers. This highlights the potential benefit of explicitly enforcing the initial condition. Additionally, for the wave equation, the prediction error for FNO and CNO increases as time evolves. In contrast, the RONOM method maintains a relatively stable error over time. This suggests that the causal structure and inductive bias introduced by the neural ODE in RONOM help mitigate error propagation.
In conclusion, while FNO outperforms CNO and RONOM on the Burgers’ equation, RONOM, which fundamentally is a classical vector-to-vector neural network, still demonstrates competitive performance on the Burgers and wave equations.
Figure 3: Input generalization (Burgers’ equation). Visualizations of FNO, CNO, DeepONet, and RONOM model predictions over time on the Burgers’ equation. The ground truth (GT) solution and the corresponding discrepancy between GT and model prediction are also demonstrated for comparison.
Figure 4: Input generalization (wave equation). Visualization of prediction errors for the FNO, CNO, DeepONet, and RONOM models, shown alongside the ground truth (GT).
4.2.2 Super-resolution
Figure 5: Super-resolution (wave equation). Visualization of prediction errors for the FNO, CNO, DeepONet, and RONOM models, shown alongside the ground truth (GT). The initial condition, discretized on the training grid, is mapped to the output at training spatial resolution and full time resolution. Time instances highlighted in bold refer to the times that were not included during training.
The performance of FNO, CNO, DeepONet, and the proposed RONOM on super-resolution is also shown in Table 1. The result in the Time column shows that the FNO still performs best on both Burgers’ equation and wave equation, while the DeepONet performs the worst. CNO and RONOM demonstrate competitive performance with the FNO on the Burgers’ equation. But the performance of both RONOM and CNO deteriorates more compared to that of the FNO on the wave equation.
Figure 5 further illustrates such deterioration by demonstrating the error of the model predictions at unseen time instances. While FNO closely matches the intermediate solutions, both the CNO and RONOM exhibit significantly higher errors at these times. A key factor that leads to such behavior is the application of global Fourier layers versus local convolutional layers. The Fourier layers allow FNO to have global temporal interactions, while CNO and RONOM rely on local operations, which might be insufficient for generalizing in time. Additionally, the inductive bias introduced by the neural ODE formulation in RONOM might be overly restrictive, resulting in poorer reconstructions at unseen time points. Another interesting observation from Table 1 is that CNO appears to struggle with spatial upsampling, despite being specifically designed for this. This issue is further discussed in the next section, which focuses on robustness with respect to input discretization.
4.3 Results on discretization robustness
This section investigates the behaviors of the FNO, CNO, and RONOM models under varying discretizations of the input and output. Since DeepONet requires inputs remaining at the training resolution, it is excluded from this comparison.
For the Burgers’ equation, spatial subsampling by factors of 64, 32, 16, 8, and 1 is performed, where a factor of 32 is used for the training. For the wave equation, spatial subsampling is done by factors of 8, 4, 2, and 1, with a factor of 4 matching the training resolution. For the time variable, performance is evaluated both at the training resolution of 11 time points and at the full resolution of 101 time points. The results of this experiment for the Burgers’ equation and wave equation are presented in Figure 6(a) and 6(b), respectively.
(a) Burgers’ equation
(b) Wave equation
Figure 6: Discretization robustness performance. The quantitative performance of the FNO, CNO, and RONOM models is compared when discretizing the inputs and outputs at different spatial resolutions, with the dashed black line indicating the training spatial resolution. The results are shown both at the time discretization used during training and at the full time resolution available in the training data. RONOM is the only method robust to spatial resolution changes.
Figure 7: Discretization effects (Burgers’ equation). The first column presents the ground truth solution at both the training resolution and the full resolution. The subsequent columns display the pixel-wise reconstruction errors for the FNO, CNO, and RONOM methods.
Only the RONOM method is robust to changes in spatial input discretization. This indicates that simple convolutional architectures, as commonly used in deep learning-based ROMs, can achieve discretization robustness when equipped with appropriate lifting operations. In Raonic et al. [39], U-Net was found to be unsuitable for dealing with different input discretizations. It is mainly because spatial convolutions were defined on the training grid. When the input resolution increases but the kernel size remains fixed, the convolution spans a smaller region in physical space. As a result, it effectively operates on smaller patches than it did during training, breaking consistency with the infinite-dimensional interpretation of the input function. By first lifting the input to a function and then querying this function at the training resolution, this inconsistency can be avoided, and it remains discretization robust.
To better understand why RONOM remains robust, while FNO and CNO do not, Figures 7 and 8 display predictions for FNO, CNO, and RONOM on the Burgers and wave equations, respectively. These figures confirm that, while the FNO achieves strong performance at the training resolution, it fails to generalize to the full spatial resolution. The CNO also exhibits poor performance at this higher resolution, primarily due to an inaccurate reconstruction of the initial condition.
Figure 8: Discretization effects (wave equation). The top row shows a ground truth example at the training resolution and the full data resolution. Below are the pixel-wise errors for the FNO, CNO, and RONOM methods.
In the architecture of CNO, inputs at resolutions different from the training resolution are resampled to the training resolution before processing. The resulting outputs are also resampled to match the desired output resolution. According to the Nyquist sampling theorem, a specific grid resolution is equivalent to a function in some space of bandlimited functions. Therefore, varying resolutions require mapping between different spaces of bandlimited functions. Specifically, consider a function $f$ whose Fourier transform satisfies $\operatorname{supp}(\hat{f})\subseteq[-\omega,\omega]^{d}$. Upsampling to a space with a broader Fourier support, i.e. $[-\underline{\omega},\underline{\omega}]^{d}$ with $\underline{\omega}\geq\omega$, corresponds to the identity operation as no new frequency components are introduced. When downsampling to a function with Fourier support in $[-\underline{\omega},\underline{\omega}]^{d}$ where $\underline{\omega}<\omega$, CNO uses $\tilde{f}(x)=\left(\underline{\omega}/{\omega}\right)^{d}\int_{\Omega}h_{\underline{\omega}}(x-y)f(y)\mathrm{d}y$ where $h_{\underline{\omega}}=\prod_{i=1}^{d}\operatorname{sinc}(2\underline{\omega}x_{i})$ is the interpolation sinc filter with Fourier transform $\hat{h}_{\underline{\omega}}=\mathbbm{1}_{[-\underline{\omega},\underline{\omega}]^{d}}$. Since convolution in the spatial domain corresponds to multiplication in the Fourier domain, this operation effectively filters out the frequency components outside the new bandlimit. In practice, the CNO performs both upsampling and downsampling in Fourier space. Downsampling is achieved by truncating high-frequency components, while upsampling involves zero-padding the Fourier coefficients. This procedure uses the Fast Fourier Transform and only works properly when the source and target grid sizes are integer multiples of one another. When this condition is not met, the original Fourier coefficients correspond to incorrect (shifted) spatial frequencies in the new truncated or zero-padded representation. As a result, a different function is represented, producing errors compared to the original function. This makes the CNO sensitive to the specific discretization of the input. In contrast, our encoder employs a projection-based approach for up- and downsampling that is independent of grid structures. It supports arbitrary input locations, as long as the approximations of the inner products in Theorem 2.3 satisfy suitable error bounds (as discussed in Section 2.3.3). This explains the robustness of our method compared to CNO and also accounts for the failure of the CNO on the spatial super-resolution experiment reported in Table 1.
5 Conclusion
ROM and neural operators are both widely used to speed up the computation of solutions to time-dependent PDEs. While ROM methods rely on a fixed input discretization, neural operators are designed to generalize across varying discretizations. However, neural operator approaches do not explicitly quantify how well the discretized operators approximate their learned infinite-dimensional counterparts. In contrast, the ROM framework explicitly accounts for the discretization error of the associated ordinary differential equations. This work bridges the concepts from ROM and operator learning by introducing the reduced-order neural operator modeling (RONOM) framework. A general error estimate is derived to quantify the discrepancy between the learned infinite-dimensional operator and its approximation, explicitly accounting for errors arising from both the input function discretization and the neural ODE discretization. The error estimate guarantees discretization convergence of our model. Furthermore, RONOM is compared to FNO, CNO, and DeepONet on two numerical examples on three key aspects: input generalization, super-resolution, and robustness to discretization changes. The results show that RONOM, which uses standard vector-to-vector neural networks, achieves competitive generalization and exhibits strong performance in spatial super-resolution and discretization robustness, where FNO and CNO underperform. Moreover, our findings on temporal super-resolution provide insights for improving temporal accuracy in operator learning.
6 Future work
To further leverage the knowledge about theory and computational aspects of ROM, it could be valuable to further explore ROM’s low-dimensional representation and the related Kolmogorov n-width. In particular, our methodology could benefit from incorporating structures into the learned latent space, like structure-preserving ROMs. These ROMs retain physical or geometric properties, such as energy conservation, symplecticity, or passivity in the reduced system to ensure stability and accuracy over long time horizons [35, 34, 9]. Physics-informed techniques [25, 47, 13] may also be useful by explicitly incorporating the PDE into the loss function. Moreover, ROM sometimes explicitly enforces the initial condition, a feature not currently implemented in our approach. More broadly, integrating additional low-dimensional, structure-preserving, and physics-informed strategies within our RONOM offers a promising direction for future work.
Furthermore, while our method currently only considers initial conditions as the input of the operator, it can be extended to other types of inputs. For example, in scenarios involving a spatially varying diffusivity function with a fixed initial condition, the encoder could map the input to a conditioning vector for the neural ODE. This would allow the neural ODE to exhibit different trajectories within the latent space while maintaining the same initial latent code, and thus, the same initial condition. Extending the RONOM framework in this direction could be a promising direction for future work.
Competing interests
All authors declare that they have no conflicts of interest.
Data availability
All the data and source codes to reproduce the results will be made available on GitHub at https://github.com/SCdummer/RONOM.git.
Appendix A Proof of Theorem 2.5
Before proving the theorem, we state and prove the following lemma
Lemma A.1.
With $\lambda$ the Lebesgue measure, assume $\Omega=\bigcup_{k=1}^{N}\Omega_{k}$ with $\lambda(\Omega_{i}\cap\Omega_{j})=0$ for $i\neq j$ and $\lambda(\Omega_{k})>0$. Given a $u\in\operatorname{TV}(\Omega)$ and denoting the restriction to $\Omega_{k}$ as $u\mid_{\Omega_{k}}\colon\Omega_{k}\to\mathbb{R}$, the following holds,
$\operatorname{TV}(u):=\sup_{\begin{subarray}{c}\phi\in C_{c}^{\infty}(\Omega)\\ \norm{\phi}_{L_{\infty}(\Omega)}\leq 1\end{subarray}}\int_{\Omega}u(\bm{x})\nabla\cdot\phi(\bm{x})\mathrm{d}\bm{x}\geq\sum_{k=1}^{N}\operatorname{TV}(u\mid_{\Omega_{k}}).$
Proof A.2.
As $\operatorname{TV}(u|_{\Omega_{k}})\leq\operatorname{TV}(u)<\infty$ due to $\Omega\subseteq\Omega_{k}$, by the definition of the total variation, we have for every $\epsilon>0$ a $\phi_{k}^{\epsilon}\in C_{c}^{\infty}(\Omega_{k})$ such that $\int_{\Omega_{k}}u(\bm{x})\nabla\cdot\phi_{k}^{\epsilon}(\bm{x})\mathrm{d}\bm{x}\geq\operatorname{TV}(u\mid_{\Omega_{k}})-\epsilon$. Define $\phi^{\epsilon}(\bm{x}):=\sum_{k=1}^{N}\phi_{k}^{\epsilon}(\bm{x})$ with $\phi_{k}^{\epsilon}(\bm{x})=0$ if $\bm{x}\in\Omega/\Omega_{k}$. Note that $\phi^{\epsilon}\in C_{c}^{\infty}(\Omega)$ and $\phi^{\epsilon}\mid_{\Omega_{k}}=\phi_{k}^{\epsilon}$ as $\lambda(\Omega_{i}\cap\Omega_{j})=0$ for $i\neq j$. Hence, we obtain:
$\operatorname{TV}(u)$ $\geq\int_{\Omega}u(\bm{x})\nabla\cdot\phi^{\epsilon}(\bm{x})\mathrm{d}\bm{x}=\sum_{k=1}^{N}\int_{\Omega}u(\bm{x})\nabla\cdot\phi_{k}^{\epsilon}(\bm{x})\mathrm{d}\bm{x}=\sum_{k=1}^{N}\int_{\Omega_{k}}u(\bm{x})\nabla\cdot\phi_{k}^{\epsilon}(\bm{x})\mathrm{d}\bm{x}$
$\geq\sum_{k=1}^{N}\operatorname{TV}(u\mid_{\Omega_{k}})-\epsilon=\left(\sum_{k=1}^{N}\operatorname{TV}(u\mid_{\Omega_{k}})\right)-\epsilon N.$
Letting $\epsilon\rightarrow 0$ yields the desired inequality.
Proof A.3 (Proof Theorem 2.5).
Note that:
(13) $\begin{split}\int_{\Omega}f(\bm{x})\mathrm{d}\bm{x}-\sum_{k=1}^{N}\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)&=\left(\int_{\Omega}f(\bm{x})\mathrm{d}\bm{x}-\int_{\Omega^{N}}f(\bm{x})\mathrm{d}\bm{x}\right)+\\ &\left(\int_{\Omega^{N}}f(\bm{x})\mathrm{d}\bm{x}-\sum_{k=1}^{N}\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right).\end{split}$
The first part of the error (domain approximation error) can be bounded as follows,
(14) $\left|\int_{\Omega}f(\bm{x})\mathrm{d}\bm{x}-\int_{\Omega^{N}}f(\bm{x})\mathrm{d}\bm{x}\right|$ $\leq\int_{\Omega\Delta\Omega^{N}}|f(\bm{x})|\mathrm{d}\bm{x}\leq|\Omega\Delta\Omega^{N}|\norm{f}_{L_{\infty}(\Omega)},$
where the final equality follows from the fact that $\operatorname{TV}$ functions are essentially bounded and $C^{p+1}(\Omega)$ functions on a compact domain are bounded. In the second part of the error, the element-wise integral for $f\in\operatorname{TV}(\Omega)$ can be decomposed into two parts,
$\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}$ $=\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}=\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}\int_{\Omega_{k}}f(\bm{x}_{i,k})+\left(f(\bm{x})-f(\bm{x}_{i,k})\right)\mathrm{d}\bm{x}$
$=|\Omega_{k}|\left(\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}f(\bm{x}_{i,k})\right)+\int_{\Omega_{k}}\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}\left(f(\bm{x})-f(\bm{x}_{i,k})\right)\mathrm{d}\bm{x}.$
As $u\in\operatorname{TV}(\Omega)$, we know $u$ is essentially bounded. Hence, there exists $c_{k}\leq f(\bm{x})\leq C_{k}$ for $\bm{x}\in\Omega_{k}$. We obtain that,
$\left|\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}-\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right|$ $=\left|\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}-|\Omega_{k}|\left(\sum_{i=1}^{m_{k}}\widetilde{w}_{i,k}f(\bm{x}_{i,k})\right)\right|$
$=\left|\sum_{i=1}^{m}\widetilde{w}_{i,k}\int_{\Omega_{k}}(f(\bm{x})-f(\bm{x}_{i,k}))\mathrm{d}\bm{x}\right|$
$\leq\sum_{i=1}^{m}\widetilde{w}_{i,k}\int_{\Omega_{k}}(C_{k}-c_{k})\mathrm{d}\bm{x}$
$=|\Omega_{k}|(C_{k}-c_{k})\leq|\Omega_{k}|\operatorname{TV}(f\mid_{\Omega_{k}}),$
where the first inequality uses $c_{k}\leq f(\bm{x})\leq C_{k}$ for $\bm{x}\in\Omega_{k}$ and the final inequality uses the definition of the total variation. Furthermore, from [14], it follows that for $f\in C^{p+1}(\Omega)$, $\left|\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}-\sum_{i=1}^{m}w_{i,k}f(\bm{x}_{i,k})\right|\leq c|\Omega_{k}|h_{\Omega_{k}}^{p+1}\sup_{\begin{subarray}{c}\bm{x}\in\Omega_{k}\\ \sum_{j=1}^{d}\gamma_{j}=p+1\end{subarray}}\left|\partial^{\gamma}f(\bm{x})\right|$. Utilizing this, the following error estimate on $\Omega^{N}$ is obtained,
$\left|\int_{\Omega^{N}}f(\bm{x})\mathrm{d}\bm{x}-\sum_{k=1}^{N}\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right|$ $=\left|\sum_{k=1}^{N}\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}-\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right|$
$\leq\sum_{k=1}^{N}\left|\int_{\Omega_{k}}f(\bm{x})\mathrm{d}\bm{x}-\left(\sum_{i=1}^{m_{k}}w_{i,k}f(\bm{x}_{i,k})\right)\right|$
$\leq\begin{cases}c\sum\limits_{k=1}^{N}h_{\Omega_{k}}^{p+1}|\Omega_{k}|\sup_{\begin{subarray}{c}\bm{x}\in\Omega_{k}\\ \sum_{i=1}^{d}\gamma_{i}=p+1\end{subarray}}\left|\partial^{\gamma}f(\bm{x})\right|\\ \sum\limits_{k=1}^{N}|\Omega_{k}|\operatorname{TV}(f\mid_{\Omega_{k}})\end{cases}$
$\leq\begin{cases}ch^{p+1}|\Omega^{N}|\sup_{\begin{subarray}{c}\bm{x}\in\Omega\\ \sum_{i=1}^{d}\gamma_{i}=p+1\end{subarray}}\left|\partial^{\gamma}f(\bm{x})\right|\\ h\operatorname{TV}(f),\end{cases}$
where the first case corresponds to $f\in C^{p+1}(\Omega)$ and the second to $f\in\operatorname{TV}(\Omega)$. The second inequality uses the error estimates on $\Omega_{k}$, and the last inequality follows from Lemma A.1 and $h_{\Omega_{k}}\leq h$. By combining the above estimate with inequalities (13) and (14), the desired error estimate is obtained.